# Lubrication-mediated rebounds off fluid baths

### v1

@@ 1 Introduction
The role of the entrained air layer present throughout droplet rebounds has been well studied. Since the work of Worthington (1881), where it was first questioned, it has been well established that rebounds are only able to occur in the presence of a lubricating layer of air preventing the two liquids to have contact. In the case of immiscible fluids, or one of the impactors being solid, the same assumption can be made in the pre-impact time; that the layer of air helps aid the rebound to occur (Moore, 2021). Recent studies of drops that bounce exist in several configurations, for example off solid hydrophobic surfaces Kolinski et al. (2014) and off liquid films Tang et al. (2019).
There are fewer studies of drops rebounding off deep baths, particularly in the regime where the underlying fluid is inviscid (or nearly so) and capillary waves are long lived after the droplet impact. These models gained interest in part because of the work of Couder on indefinitely bouncing drops on vibrating baths Couder et al. (2005) and its fascinating quantum mechanical analogies (Bush (2015) and references therein). The studies of (Galeano-Rios et al., 2017, 2019) and Alventosa et al. (2023) imposed a so-called kinematic match between the solid sphere or drop and the bath to avoid modelling the air layer and obtained accurate results when compared to Faraday-pilot wave studies of bouncing and walking droplets. In (Galeano-Rios et al., 2021) the kinematic match model was validated against experiments and DNS simulations of the Navier-Stokes equations of single rebounds of solid hydrophobic spheres for low Weber number impacts.
@@ 2 Model
$p_{\beta}\mathbf{n}_{\beta}+\mu_{\beta}\tau_{\beta}\mathbf{n}_{\beta}$ $=p_{a}\mathbf{n}+\mu_{a}\tau_{a}\mathbf{n}_{\beta}-\sigma_{\beta}\kappa_{\beta},\qquad$ $\mathbf{x}$ $\in\partial\Omega_{\beta}.$ (4)
Where $\tau_{\alpha}=(\nabla\mathbf{u}_{\alpha}+\nabla\mathbf{u}_{\alpha}^{T})$ is the stress tensor for each fluid, $\kappa_{\beta}$ denotes curvature of the free surface with outward unit normal vector $\mathbf{n}_{\beta}$, and $D_{t}$ is the material derivative. These equations are supplemented by imposing decay of velocities in the far field or appropriate boundary conditions in the case of finite domains. We proceed with several modelling approximations which will reduce the system to a tractable state.
@@ 2.1 Linear quasi-potential model for the liquid bath
We assume viscous effects are small except in the lubrication layer, and follow methodology from Dias et al. (2008), and take a Helmholtz decomposition of the liquid as small perturbation from potential flow $\mathbf{u}_{l}=\nabla\phi+\nabla\times\psi$. The $\psi$ terms are then written in terms of $\phi$ through a boundary layer argument. The free surface of the bath is given by $z=\eta_{b}(\mathbf{x},t)$ where $z=0$ is its undisturbed free surface. The governing linear system then becomes
00 0 $=\Delta_{H}\phi+\partial_{z}^{2}\phi,\qquad$ $z$ $\leq 0,$ (5)
$\partial_{t}\phi$ $=-\frac{1}{\rho_{b}}p_{a}+g\eta_{b}+\frac{\sigma}{\rho_{b}}\Delta_{H}\eta_{b}+2\nu_{l}\Delta_{H}\phi,\qquad$ $z$ $=0,$ (6)
$\partial_{t}\eta_{b}$ $=\partial_{z}\phi+2\nu_{l}\Delta_{H}\eta,\qquad$ $z$ $=0,$ (7)
where $\Delta_{H}$ is used to denote the horizontal Laplacian, and velocity continuity $\mathbf{u}_{b}=\mathbf{u}_{a}$ is imposed at the interface $\eta_{b}$ between the bath and air layer.
@@ 2.2 Air layer lubrication approximation
The lower part of the droplet is described by $z=\eta_{d}(\mathbf{x}_{H},t)=Z(t)-S(\mathbf{x}_{H},t)$, composed of the shape of the droplet $S$, and the vertical height of its centre of mass $Z$. The domain of $S$ in $\mathbf{x}_{H}$ is chosen such that $\nabla_{H}$ is well defined. The dynamics of the air layer are described through a lubrication approximation as in Phillips et al. (2024). Consider a lubrication region $\Omega_{a}=\Omega_{L}\times[\eta_{b}(r,\theta,t),\eta_{d}(r,\theta,t)]$ where its footprint is described in cylindrical coordinates as
@@
where $r^{*}$ is the edge of the lubrication region. Denoting $h=\eta_{d}-\eta_{b}$, the criterion for the edge of the lubrication layer is given by a criterion $h(r^{*},\theta,t)=\varepsilon$. Assuming that the layer between the droplet and the bath is thin, the leading order balance of the Navier-Stokes equations result in the lubrication balance where $\partial_{z}p_{a}=0$ with
@@
$\mu\partial_{z}^{2}\mathbf{u}_{a}^{H}$ $=\nabla_{H}p_{a},\qquad$ $\mathbf{x}$ $\in\Omega_{a},$ (10)
@@ 2.3.1 Droplet Deformation
Linear damped droplet oscillations are well understood (Lamb (1924)) and the results have been used in a variety of contexts e.g. (Aalilija et al., 2020). In this section we write equivalent equations for these oscillations using the same quasi-potential theory used above for the bath. Consider a capillary scale droplet of liquid which when unperturbed is a sphere of radius $R_{0}$. Describing the problem using spherical polar coordinates with azimuthal angle $\theta$ and polar angle $\varphi$, we can take a Helmholtz decomposition for the liquid velocity $\mathbf{u}_{d}=\nabla\phi_{d}+\nabla^{\perp}\psi$, and using a boundary layer argument (Phillips et al. (2024)) obtain similar governing equations to the bath equations (5)-(7)
@@
$\partial_{t}\phi_{d}$ $=-\frac{1}{\rho_{d}}p_{a}+\frac{\sigma}{\rho_{d}}\kappa_{d}+2\nu_{d}\partial_{r}^{2}\phi_{d},$ $r=R_{0},$ (16)
@@
$\kappa_{d}=-\left(\frac{2}{R_{0}^{2}}\zeta_{d}+\frac{1}{R_{0}^{2}}\Delta_{s}\zeta_{d}\right),\qquad\Delta=\frac{1}{r^{2}}\partial_{r}(r^{2}\partial_{r})+\frac{1}{r^{2}}\Delta_{s},$
with $\Delta_{s}$ the Laplace-Beltrami operator (i.e. the surface Laplacian). The equations are valid in the small oscillation limit and have been truncated to contain terms that are linear in the Ohnesorge number $Oh=\frac{\mu}{\sqrt{\rho R_{0}\sigma}}$. One may proceed with an eigenfunction decomposition of $\zeta_{d}$ and $\phi_{d}$ as
$\phi_{d}=\sum_{l=2}^{\infty}\sum_{m=-l}^{l}a_{lm}(t)r^{l}Y_{l}^{m},\qquad\zeta_{d}=\sum_{l=2}^{\infty}\sum_{m=-l}^{l}c_{lm}(t)Y_{l}^{m},$
@@
where $\hat{p}_{l,m}$ is the $Y_{l,m}$ coefficient of of the lubrication air pressure $p_{a}$. The damping coefficients and modal frequencies are given by
@@
as seen in the literature (Lamb, 1924; Aalilija et al., 2020; Alventosa et al., 2023).
@@ 2.3.2 Equations for the centre of mass
where $\mathbf{n_{d}}$ is the outward normal of the droplet $m=4\rho\pi R_{0}^{3}/3$ is its mass, $g$ is the gravitational constant and $\hat{\mathbf{z}}$ is the unit vector in the vertical direction. The boundary of the droplet is now given by $\mathbf{X}+(R_{0}+\eta_{d}(\theta,\varphi))\mathbf{\hat{r}}$, where $\mathbf{\hat{r}}$ is a radial unit vector relative to the position $\mathbf{X}$.
@@ 2.4 The lubrication-mediated droplet-bath system
$\partial_{t}\Phi_{b}$ $=-\frac{1}{\rho_{b}}p_{a}+\frac{\sigma}{\rho_{b}}\Delta_{H}\eta_{b}+2\nu_{l}\Delta_{H}\Phi_{b},$ (21)
$\partial_{t}\eta_{b}$ $=D_{b}\Phi_{b}+2\nu_{l}\Delta_{H}\eta_{b}\doteq F_{b},$ (22)
$\partial_{t}\Phi_{d}$ $=-\frac{1}{\rho_{d}}p_{a}+\frac{\sigma}{\rho_{d}}\kappa_{d}+2\nu_{d}\partial_{r}^{2}\Phi_{d},$ (23)
$\partial_{t}\eta_{d}$ $=D_{d}\Phi_{d}+2\nu_{d}\left(\frac{1}{r^{2}}\Delta_{s}\eta_{d}-\frac{1}{r^{3}}\Delta_{s}D_{d}^{-1}\Phi_{d}\right)\doteq F_{d},$ (24)
@@
In the surface potential equations, the pressure $p_{a}$ is nonzero only in $\Omega_{L}$ and $W=\frac{d\mathbf{X}}{dt}$ is the vertical velocity of the droplet’s centre of mass. Equation (26), where $Q$ is given by (13) is an inhomogeneous elliptic equation for the pressure with the boundary condition that $p_{a}=0$ on the free boundary $\partial\Omega_{L}$.
@@ 3 Axisymmetric solid impacts and numerical results
We now consider the simplest case in which this framework can be used and compared to prior simulations and experiments: the rebound of solid (hydrophobic) spheres Galeano-Rios et al. (2021). This simplifies considerably the system, reducing the dimension and eliminating droplet oscillations. Considering the axisymmetric domain for the bath to of finite radius $r=L$, and imposing Neumann boundary conditions there for both $\eta_{b}$ and $\Phi_{b}$ we expand these in terms of Bessel functions of the first kind,
@@
where the $k_{j}$ satisfy $J_{0}^{\prime}(k_{j}L)=0$. We note that in this basis $\phi_{b}=\sum_{j=1}^{\infty}b_{j}J_{0}(k_{j}r)e^{k_{j}z}$ to satisfy Laplace’s equation and hence the DtN map is expressed simply as
$D_{b}\Phi=\partial_{z}\phi_{b}|_{z=0}=\sum_{j=1}^{\infty}k_{j}b_{j}J_{0}(k_{j}r).$
@@
$\dot{b}_{j}$ $=-\frac{1}{\rho_{l}}\hat{p}_{j}-\frac{\sigma}{\rho_{l}}k^{2}a_{j}-2\nu_{l}k^{2}b_{j},$ (29)
@@
where $W=\dot{Z}$, $h=Z-\sqrt{R_{0}^{2}-r^{2}}-\eta_{b}$ for $r\leq R_{0}$, $r^{*}$ is defined by $h(r^{*},t)=\epsilon$ and where $\hat{p}_{j}$ are the Bessel coefficients of the Pressure provided by
@@
$p_{a}=-\int_{r^{*}}^{r}\left[\frac{12\mu_{a}}{r^{\prime}h(r^{\prime},t)^{3}}\left(\int_{0}^{r^{\prime}}r^{\prime\prime}\partial_{t}h(r^{\prime\prime},t)\text{d}r^{\prime\prime}\right)+\frac{6\mu_{a}}{h(r^{\prime},t)^{2}}\partial_{r}\Phi_{b}(r^{\prime},t)\right]\text{d}r^{\prime}\;.$
@@ 3.1 Numerical implementation
We shall compute a solutions approximating the system (28)-(3) by truncating the Bessel expansions at a large $N$ and integrating the resulting differential equations (28)-(30) with a fourth order Runge-Kutta scheme. The equation for the pressure (3) is calculated using a trapezium rule. Since Bessel expansions are ill-conditioned on regular grids, we use an oversampled nonuniform grid in $r$ with $M$ points distributed on Chebyshev collocation points on $[0,L]$ using $\theta_{j}=(j-1)\pi/(M-1)$ with $j=1\ldots M$ and $r_{j}=\frac{L(1+\cos(\theta_{j}))}{2}$. Hence we compute two matrices: an $M\times N$ matrix which evaluates an N-term Bessel series at $r_{j}$ and an $N\times M$ matrix corresponding to calculating the projection formula (32) for the Bessel coefficients of a function given at $r_{j}$.

### v2

@@ 1 Introduction
The role of the entrained air layer present throughout droplet rebounds has been well studied. Since the work of Worthington (1881), where it was first questioned, it has been well established that rebounds are only able to occur in the presence of a lubricating layer of air preventing the two liquids contacting. The assumption that within the near-contact region, the air can behave like a lubrication layer (Hicks & Purvis, 2010) has directed a reach field of literature of impacts on other substrates such as porous media (Hicks & Purvis, 2017). In the case of the impactor being solid, the same assumption can be made in the pre-impact time; that the layer of air helps delay contact (Moore, 2021). Recent studies of drops that bounce exist in several configurations, for example off solid hydrophobic surfaces (Kolinski et al., 2014), and off liquid films (Tang et al., 2019).
There are fewer studies of drops rebounding off deep baths, particularly in the regime where the underlying fluid is inviscid (or nearly so) and capillary waves are long lived after the droplet impact. These models gained interest in part because of the work of Couder on indefinitely bouncing drops on vibrating baths (Couder et al., 2005) and its fascinating quantum mechanical analogies (Bush (2015) and references therein). The studies of Galeano-Rios et al. (2017, 2019) and Alventosa et al. (2023) imposed a so-called kinematic match between the solid sphere or drop and the bath to avoid modelling the air layer and obtained accurate results when compared to Faraday-pilot wave studies of bouncing and walking droplets. In Galeano-Rios et al. (2021) the kinematic match model was validated against experiments and DNS simulations of the Navier-Stokes equations of single rebounds of solid hydrophobic spheres for low Weber number impacts.
@@ 2 Model
$p_{\beta}\mathbf{n}_{\beta}+\mu_{\beta}\tau_{\beta}\mathbf{n}_{\beta}$ $=p_{a}\mathbf{n}_{\beta}+\mu_{a}\tau_{a}\mathbf{n}_{\beta}-\sigma_{\beta}\kappa_{\beta},\qquad$ $\mathbf{x}$ $\in\partial\Omega_{\beta}.$ (4)
Where $\tau_{\alpha}=(\nabla\mathbf{u}_{\alpha}+\nabla\mathbf{u}_{\alpha}^{T})$ is the stress tensor for each fluid, $\kappa_{\beta}$ denotes curvature of the free surface with outward unit normal vector $\mathbf{n}_{\beta}$, gravity $g$ acting downwards in the $-\mathbf{e}_{z}$ direction, and $D_{t}$ is the material derivative. These equations are supplemented by imposing decay of velocities in the far field or appropriate boundary conditions in the case of finite domains. We proceed with several modelling approximations which will reduce the system to a tractable state.
@@ 2.1 Linear quasi-potential model for the liquid bath
We assume viscous effects are small except in the lubrication layer, and follow methodology from Dias et al. (2008) which we briefly outline below. Take a Helmholtz decomposition of the liquid as small perturbation from potential flow $\mathbf{u}_{b}=\nabla\phi_{b}+\nabla\times\boldsymbol{\psi_{b}}$, and use tangential stress conditions and a boundary layer argument to eliminate the $\boldsymbol{\psi}_{b}$ terms, expressing the rotational part of the flow in terms of $\phi_{b}$ and $\eta_{b}$. A more detailed explanation of this argument can be found in Phillips et al. (2024). The free surface of the bath is given by $z=\eta_{b}(\mathbf{x},t)$ where $z=0$ is its undisturbed free surface. The governing linear system then becomes
00 0 $=\Delta_{H}\phi_{b}+\partial_{z}^{2}\phi_{b},\qquad$ $z$ $\leq 0,$ (5)
$\partial_{t}\phi_{b}$ $=-\frac{1}{\rho_{b}}p_{a}-g\eta_{b}+\frac{\sigma_{b}}{\rho_{b}}\Delta_{H}\eta_{b}+2\nu_{b}\Delta_{H}\phi_{b},\qquad$ $z$ $=0,$ (6)
$\partial_{t}\eta_{b}$ $=\partial_{z}\phi_{b}+2\nu_{b}\Delta_{H}\eta_{b},\qquad$ $z$ $=0,$ (7)
where $\Delta_{H}$ is used to denote the horizontal Laplacian, the curvature is small and of the form $\kappa_{b}=-\nabla_{H}\eta_{b}$, and velocity continuity $\mathbf{u}_{b}=\mathbf{u}_{a}$ is imposed at the interface $\eta_{b}$ between the bath and air layer. The terms proportional to $\nu_{b}$ arise from $\boldsymbol{\psi_{b}}$, and are the corrections to the pressure and vertical velocity, respectively.
@@ 2.2 Air layer lubrication approximation
Away from the droplet-bath interaction, the air is assumed to be atmospheric pressure and have negligible effect on the system, as we must introduce the domain in which the lubrication effects may be important. The lower part of the droplet is described by $z=\eta_{d}(\mathbf{x}_{H},t)=\mathbf{X}(t)-S(\mathbf{x}_{H},t)$, composed of the shape of the lower part of the droplet $S$, and the vertical height of its centre of mass $\mathbf{X}$. The domain of $S$ in $\mathbf{x}_{H}$ is chosen such that $S$ is single valued and with bounded gradients. The dynamics of the air layer are described through a lubrication approximation as in Phillips et al. (2024). Consider a lubrication region $\Omega_{a}=\Omega_{L}\times[\eta_{b}(r,\theta,t),\eta_{d}(r,\theta,t)]$ where its footprint is described in cylindrical coordinates as
@@
where $r^{*}$ is the edge of the lubrication region. Denoting $h=\eta_{d}-\eta_{b}$, the edge of the lubrication layer is given by a criterion $h(r^{*},\theta,t)=\varepsilon$. It is assumed that the $\Omega_{L}$ is a subset of the domain of S. Assuming that the layer between the droplet and the bath is thin, the leading order balance of the Navier-Stokes equations result in the lubrication balance where $\partial_{z}p_{a}=0$ with
@@
$\mu_{a}\partial_{z}^{2}\mathbf{u}_{a}^{H}$ $=\nabla_{H}p_{a},\qquad$ $\mathbf{x}$ $\in\Omega_{a},$ (10)
@@ 2.3.1 Droplet Deformation
Linear damped droplet oscillations are well understood (Lamb, 1924) and the results have been used in a variety of contexts e.g. Aalilija et al. (2020). In this section we write equivalent equations for these oscillations using the same quasi-potential theory used above for the bath. Consider a capillary scale droplet of liquid which when unperturbed is a sphere of radius $R_{0}$. Describing the problem using spherical polar coordinates with azimuthal angle $\theta$ and polar angle $\varphi$, we can take a Helmholtz decomposition for the liquid velocity $\mathbf{u}_{d}=\nabla\phi_{d}+\nabla\times\boldsymbol{\psi}_{d}$. Using the same argument argument as above (Phillips et al., 2024) we obtain similar governing equations to those of the bath (5)-(7)
@@
$\partial_{t}\phi_{d}$ $=-\frac{1}{\rho_{d}}p_{a}-\frac{\sigma_{d}}{\rho_{d}}\kappa_{d}-2\nu_{d}\partial_{r}^{2}\phi_{d},$ $r=R_{0},$ (16)
@@
$\kappa_{d}=-\left(\frac{2}{R_{0}^{2}}\zeta_{d}+\frac{1}{R_{0}^{2}}\Delta_{s}\zeta_{d}\right),\qquad\Delta=\frac{1}{r^{2}}\partial_{r}(r^{2}\partial_{r})+\frac{1}{r^{2}}\Delta_{s},$ (18)
with $\Delta_{s}$ the Laplace-Beltrami operator (i.e. the surface Laplacian). The equations are valid in the small oscillation limit and have been truncated to contain terms that are linear in the Ohnesorge number (see appendix for a discussion of the nondimensional equations). One may proceed with an eigenfunction decomposition of $\zeta_{d}$ and $\phi_{d}$ as
$\phi_{d}=\sum_{l=2}^{\infty}\sum_{m=-l}^{l}a_{lm}(t)r^{l}Y_{l}^{m},\qquad\zeta_{d}=\sum_{l=2}^{\infty}\sum_{m=-l}^{l}c_{lm}(t)Y_{l}^{m},$ (19)
@@
where $\hat{p}_{l,m}$ is the $Y_{l,m}$ coefficient of of the lubrication air pressure $p_{a}$. The coefficients of the velocity potential are given to leading order by $a_{l,m}=\frac{\dot{c}_{l,m}}{lR_{0}^{l-1}}$. The damping coefficients and modal frequencies are given by
@@
as seen in the literature (Lamb, 1924; Aalilija et al., 2020; Alventosa et al., 2023). A similar equation may be obtained for the coefficients $a_{lm}(t)$, however these are uneccessary for the LM model and have been omitted for brevity.
@@ 2.3.2 Equations for the centre of mass
where $\mathbf{n_{d}}$ is the outward normal of the droplet $m=4\rho\pi R_{0}^{3}/3$ is its mass, $g$ is the gravitational constant and $\hat{\mathbf{z}}$ is the unit vector in the vertical direction. The boundary of the droplet is now given by $\mathbf{X}+(R_{0}+\zeta_{d}(\theta,\varphi))\mathbf{\hat{r}}$, where $\mathbf{\hat{r}}$ is a radial unit vector relative to the position $\mathbf{X}$.
@@ 2.4 The lubrication-mediated droplet-bath system
$\partial_{t}\Phi_{b}$ $=-\frac{1}{\rho_{b}}p_{a}-g\eta_{b}+\frac{\sigma_{b}}{\rho_{b}}\Delta_{H}\eta_{b}+2\nu_{b}\Delta_{H}\Phi_{b},$ (24)
$\partial_{t}\eta_{b}$ $=D_{b}\Phi_{b}+2\nu_{b}\Delta_{H}\eta_{b}\doteq F_{b},$ (25)
$\partial_{t}\Phi_{d}$ $=-\frac{1}{\rho_{d}}p_{a}-\frac{\sigma_{d}}{\rho_{d}}\kappa_{d}-2\nu_{d}\partial_{r}^{2}\Phi_{d},$ (26)
$\partial_{t}\zeta_{d}$ $=D_{d}\Phi_{d}+2\nu_{d}\left(\frac{1}{r^{2}}\Delta_{s}\zeta_{d}-\frac{1}{r^{3}}\Delta_{s}D_{d}^{-1}\zeta_{d}\right)\doteq F_{d},$ (27)
@@
In the surface potential equations, the pressure $p_{a}$ is nonzero only in $\Omega_{L}$ and $W=\frac{d\mathbf{X}}{dt}$ is the vertical velocity of the droplet’s centre of mass. In the kinematic condition for the droplet we used the leading order balance $\partial_{t}\zeta_{d}=D_{d}\Phi_{d}$ to express the right hand side locally in time. While $D_{d}$ was defined using the velocity potential, its inverse acts on $\zeta_{d}$ here. This is admissible since it is an invertible linear operator on functions with mean zero, and we may restrict the inverse map to have mean zero also. Equation (29), where $\mathbf{Q}$ is given by (13) is an inhomogeneous elliptic equation for the pressure with the boundary condition that $p_{a}=0$ on the free boundary $\partial\Omega_{L}$.
@@ 3 Axisymmetric solid impacts and numerical results
We now consider the simplest case in which this framework can be used and compared to prior simulations and experiments: the rebound of solid (hydrophobic) spheres (Galeano-Rios et al., 2021). This simplifies considerably the system, reducing the dimension and eliminating droplet oscillations. Considering the axisymmetric domain for the bath to of finite radius $r=L$, and imposing Neumann boundary conditions there for both $\eta_{b}$ and $\Phi_{b}$ we expand these in terms of Bessel functions of the first kind,
@@
where the $k_{j}$ satisfy $J_{0}^{\prime}(k_{j}L)=0$. We note that in this basis $\phi_{b}=\sum_{j=1}^{\infty}b_{j}J_{0}(k_{j}r)e^{k_{j}z}$ to satisfy Laplace’s equation and the DtN map is expressed simply as
$D_{b}\Phi=\partial_{z}\phi_{b}|_{z=0}=\sum_{j=1}^{\infty}k_{j}b_{j}J_{0}(k_{j}r).$ (31)
@@
$\dot{b}_{j}$ $=-\frac{1}{\rho_{b}}\hat{p}_{j}-\frac{\sigma_{b}}{\rho_{b}}k_{j}^{2}a_{j}-2\nu_{l}k^{2}b_{j},$ (33)
@@
where $W=\dot{Z}$, $h=Z-\sqrt{R_{0}^{2}-r^{2}}-\eta_{b}$ for $r\leq R_{0}$, $r^{*}$ is defined by $h(r^{*},t)=\varepsilon$ for $\varepsilon\ll 1$, and where $\hat{p}_{j}$ are the Bessel coefficients of the Pressure provided by
@@
$p_{a}=-\int_{r^{*}}^{r}\left[\frac{12\mu_{a}}{r^{\prime}h(r^{\prime},t)^{3}}\left(\int_{0}^{r^{\prime}}r^{\prime\prime}\partial_{t}h(r^{\prime\prime},t)\text{d}r^{\prime\prime}\right)+\frac{6\mu_{a}}{h(r^{\prime},t)^{2}}\partial_{r}\Phi_{b}(r^{\prime},t)\right]\text{d}r^{\prime}\;.$ (37)
@@ 3.1 Numerical implementation
We shall compute a solution approximating the system (32)-(37) by truncating the Bessel expansions at a large $N$ and integrating the resulting differential equations (32)-(34) with a fourth order Runge-Kutta scheme. The equation for the pressure (37) is calculated using a trapezium rule. Since Bessel expansions are ill-conditioned on regular grids, we use an oversampled nonuniform grid in $r$ with $M$ points distributed on Chebyshev collocation points on $[0,L]$ using $\theta_{j}=(j-1)\pi/(M-1)$ with $j=1\ldots M$ and $r_{j}=\frac{L(1+\cos(\theta_{j}))}{2}$. Hence we compute two matrices: an $M\times N$ matrix which evaluates an N-term Bessel series at $r_{j}$ and an $N\times M$ matrix corresponding to calculating the projection formula (36) for the Bessel coefficients of a function given at $r_{j}$.